OpenStudy (anonymous):

Factor: x2 - 9x + 12

OpenStudy (unklerhaukus):

One technique that always works is to use the quadratic formula. A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Factored for of the Quadratic Equation\[(x-x_1)(x-x_2)=0\] start by identifying the coefficients a,b,c Compare \[x^2 - 9x + 12=0\] With \[ax^2+bx+c=0\] \(a=\)? \(b=\)? \(c=\)?